OpenStudy (anonymous):

Devin wants to add a street lamp to his model railroad layout. How tall should the model be if the real street lamp is 315 inches and the scale factor is 1:90?

OpenStudy (anonymous):

same idea as before x/315 = 1/90

OpenStudy (anonymous):

(1/90)(315) = x

OpenStudy (anonymous):

x=3.5

OpenStudy (anonymous):

im still lost